6x-12y=24
-x-6y=4

Found 2 solutions by CubeyThePenguin, Edwin McCravy:
Answer by CubeyThePenguin(3113)

6x-12y = 24 (eqn 1)
-x-6y = 4 (eqn 2)

Multiply equation 2 by -2:

6x - 12y =24
2x +12y = -8

Add the equations to get 8x = 16. So, (x, y) = (2,-1).


Answer by Edwin McCravy(20056)

Instead of doing it for you, I'll do one exactly like it,
step-by-step.  Just follow it as a guide.

I'll do this one instead:

7x-13y = 73
-x-7y=25

Solve the second equation for x

-x-7y=25
   -x=25+7y
    x=-25-7y   

Substitute (-25-7y) for x in the first equation

7x-13y = 73

7(-25-7y)-13y = 73

 -175-49y-13y = 73
     -175-62y = 73
         -62y = 248
            y = 248/(-62)
            y = -4

Substitute (-4) for y in

    x=-25-7y
    x=-25-7(-4)
    x =-25+28
    x = 3

Now do yours step by step the same way.

Edwin